OpenStudy (anonymous):

Abdul's Coffee Shop makes a blend that is a mixture of two types of coffee. Type A coffee costs Abdul $5.90 per pound, and type B coffee costs $4.65 per pound. This month, Abdul made $178 pounds of the blend, for a total cost of $951.45 How many pounds of type B coffee did he use?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

$178 pounds???? or should it be 178 pounds...

OpenStudy (anonymous):

yes you right

OpenStudy (anonymous):

178 pounds

OpenStudy (anonymous):

ok ok

OpenStudy (anonymous):

first let x be the total number of pounds of type A and y be the total number of pounds of type B. Now the given said that the total pound of the coffee is 178 pounds which means that 178 pounds is just the sum of the total number of type A and type B pounds which can be written as, x+y=178 <--------(1) now we know that the total cost for type A is 5.90x and the total cost for type B is 4.65y. Now the total cost for all coffee type is 951.45 dollars so 5.90x+4.65y=951.45 <-------(2) now we solve for x... from the equation one we can see that y=178-x substituting this to our equation 2 we'll have, 5.90x+4.65(178-x)=951.45 5.90x+827.7-4.65x=951.45 1.25x=951.45-827.7 1.25x/1.25=123.75/1.25 x=99 pounds <answer

OpenStudy (anonymous):

the answer is incorrect

OpenStudy (anonymous):

ah sorry... that's for type A for type B it's y=178-x y=178-99 y=79 pounds